# Detect All-Type Deepfake Audio: Wavelet Prompt Tuning for Enhanced Auditory Perception

### v1

@@
Abstract.
The rapid advancement of audio generation technologies has escalated the risks of malicious deepfake audio across speech, sound, singing voice, and music, threatening multimedia security and trust. While existing countermeasures (CMs) perform well in single-type audio deepfake detection (ADD), their performance declines in cross-type scenarios. This paper is dedicated to studying the all-type ADD task. We are the first to comprehensively establish an all-type ADD benchmark to evaluate current CMs, incorporating cross-type deepfake detection across speech, sound, singing voice, and music. Then, we introduce the prompt tuning self-supervised learning (PT-SSL) training paradigm, which optimizes SSL front-end by learning specialized prompt tokens for ADD, requiring 458× fewer trainable parameters than fine-tuning (FT). Considering the auditory perception of different audio types, we propose the wavelet prompt tuning (WPT)-SSL method to capture type-invariant auditory deepfake information from the frequency domain without requiring additional training parameters, thereby enhancing performance over FT in the all-type ADD task. To achieve an universally CM, we utilize all types of deepfake audio for co-training. Experimental results demonstrate that WPT-XLSR-AASIST achieved the best performance, with an average EER of 3.58% across all evaluation sets. The code is online available  [^1] https://github.com/xieyuankun/All-Type-ADD.
Audio Deepfake Detection, Countermeasures, Prompt Tuning, Discrete Wavelet Transform
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
With the development of audio language model (ALM) technology, it has become increasingly easy to synthesize any type of audio, including deepfake speech, sound, singing voice, and music. These deepfake audios pose a threat to society in various fields such as media, entertainment, cybersecurity, and political communication. Fortunately, research on audio deepfake detection (ADD) has been increasing annually. Among these, the earliest studies focused on deepfake speech detection. Researchers have developed a series of deepfake countermeasures (CMs) aimed at effectively detecting deepfake speech, based on the ASVspoof challenges [1, 2, 3]. Currently, some ADD research has gone beyond speech, such as the detection of deepfake singing voices [4, 5, 6], sounds [7, 8], and music [9].
Figure 1. The challenge for current single-type trained CMs toward cross-type ADD task, highlighting the effectiveness of our proposed WPT-SSL CM.
@@ 1. Introduction
For CMs, the most effective approach currently is to use pre-trained self-supervised learning (SSL) features along with a classification backbone. A representative CM in speech deepfake detection is XLSR-AASIST [10], which fine-tunes (FT) the wav2vec2-xls-r (XLSR) [11] model on speech deepfake detection dataset, achieving excellent intra-domain (ID) and out-of-domain (OOD) generalization performance. However, when dealing with the all-type ADD task, several challenges are encountered. Firstly, from the data perspective, it is uncertain whether a CM trained on single audio type can generalize to detect other types of deepfake audio. Although some studies have investigated cross-type detection for two types [12, 8], there has been no exploration of cross-type detection for all audio types. Secondly, there has been no investigation into whether a domain-invariant feature exists that can ensure the invariance of authenticity discrimination across different audio types. This requires a detailed investigation of various SSL features as well as handcrafted features. Lastly, concerning the algorithm, although fine-tuning can yield promising results, it is highly dependent on specific hyper-parameters and requires a significant amount of training parameters [13].
@@
To address the aforementioned challenges, in this paper, we aim to develop an all-type audio deepfake CM. We are the first to comprehensively establish an all-type ADD benchmark, which includes cross-type deepfake detection among speech, sound, singing voice, and music. For the feature of CMs, we investigate handcrafted features, raw waveforms, and various SSL-based features through both freezing and fine-tuning. For the classifier, we use AASIST [14], the most popular model in the field of ADD, as the back-end, and combine it with SSL front-end to form SSL-AASIST.
@@
To efficiently optimize SSL front-end, inspired by Visual Prompt Tuning (VPT) [15], we proposed the Prompt Tuning (PT)-SSL training paradigm for ADD task. PT-SSL introduces learnable prompt tokens before the input of each transformer layer, while keeping the other parameters of the layers frozen, with the goal of learning specialized prompt tokens for the ADD task. Furthermore, considering the human cognition of different audio types, the primary differences in perceiving audio types lie in their frequency domain distributions [16, 17, 18]. However, current SSL models like wav2vec2, which are primarily designed for speech recognition, focus on temporal and specific speech frequency information, lacking the ability to capture full-frequency information. To enhance frequency domain adaptability and enable SSL-based CM to quickly adapt to all types of deepfake audio, we propose wavelet prompt learning (WPT)-SSL method. WPT-SSL applies a discrete wavelet transform (DWT) to a portion of the prompt tokens, obtaining tokens for different frequency bands, thereby enhancing the full-frequency perception capability of SSL-based CM. Surprisingly, we discovered that WPT-SSL can learn a type-invariant deepfake detection prompt in a specific frequency band (HH) obtained through wavelet decomposition, thereby enabling all-type audio deepfake detection.
@@
2. Related Work
In this section, we primarily review the current popular ADD algorithms and the prompt tuning-related algorithms.
Figure 2. Our proposed PT-SSL-AASIST (left) and WPT-SSL-AASIST (right). The differences between PT and WPT are illustrated below.
2.1. Audio Deepfake Detection
In introducing the current work in the field of ADD, we categorize it based on types: speech, sound, singing voice, and music.
Speech: Speech deepfake detection is the most extensively researched and thoroughly studied task among the four types. Surrounding the ASVspoof series, a number of outstanding studies have been developed. For CM without an SSL front-end, Jung et al. [14] proposed AASIST, which achieved an EER of 0.83% on the ASVspoof2019LA (19LA) test set for the first time without SSL features, by employing raw waveform input and a spectro-temporal graph attention method. For CM with SSL front-end, Tak et al. [10] proposed using XLSR as a front-end for fine-tuning, followed by the AASIST back-end, achieving milestone scores in the ADD field on the ASVspoof2021LA (21LA) and ASVspoof2021DF (21DF) datasets. Subsequent research has focused on exploring different categories of SSL [19, 20], the utilization of SSL layers [21, 13, 22], and robustness of CM [23, 24].
Other types (sound, singing, music): Compared to the work on speech deepfake detection, research on other types is still primarily at the dataset stage, with fewer methodological explorations. For sound, Xie et al. [8] introduced the A3 subset of the Codecfake dataset, which includes deepfake sounds generated by Audiogen [25]. Furthermore, Xie et al. introduced the FakeSound dataset [7], developing novel countermeasures for sound deepfake detection. In the realm of singing voice, numerous advanced studies [26, 27, 28] have been proposed in relation to the SVDD challenge [4]. These studies indicate that, in addition to XLSR, the integration of other SSL models such as MERT [29] and WavLM [30] can further enhance CM performance. In the field of music, Luca et al. proposed FakeMusicCaps [9], which covers various ALM music synthesis methods. Concurrently, there have been limit studies on fake music detection [31, 32].
Cross types: Regarding cross-type ADD task, some studies have explored the deepfake detection effects between two audio types. For instance, Gohari et al. [12] investigated cross-type ADD between speech and singing voice, while Li et al. [33] investigated the extension of speech deepfake detection to music. However, there is currently no research covering all-audio types.
2.2. Prompt Tuning
VPT [15] was the first to successfully introduce the PT [34] method into the visual domain. It consists of the VPT-SHALLOW version, which incorporates learnable prompts before the first Transformer layer of VIT, and the VPT-DEEP version, which prepends learnable parameters to the input of each Transformer encoder layer. In VTAB-1k [35], VPT-DEEP surpasses most fine-tuning methods and conserves a substantial number of parameters. The VPT approach and its enhancements have shown promising results in multiple downstream tasks [36, 37, 38].
Owing to the strong adaptability VPT to downstream tasks and minimal storage parameters, audio researchers have begun to apply the VPT paradigm to audio tasks [39, 40]. Recently, Oiso et al. [41] investigated the use of the PT paradigm for ADD tasks with limited data, and their study was the first to demonstrate the effectiveness of the PT paradigm in the ADD domain. However, they only utilized the SHALLOW PT paradigm with the first layer combined with prompts and tuned using target domain data. The potential of PT paradigm in speech deepfake detection and all-type deepfake detection has yet to be fully explored.
3. Prompt Tuning Countermeasure
@@ 3.1. PT-SSL-AASIST
(2) $[Z_{i},E_{i}]={\color[rgb]{0,0,1}L_{i}}([{\color[rgb]{1,0,0}P_{i}},E_{i-1}]),\quad\text{for }i=2,3,...,l.$
Taking the most commonly used SSL feature in the ADD domain, XLSR-300m [^2] https://huggingface.co/facebook/wav2vec2-xls-r-300m, as an example, after the final 24 layers, we obtain a matrix output $I=[Z_{24},E_{24}]$. $I$ will serve as the input to AASIST. The back-end AASIST classifier fully follows the structure of SSL-AASIST by Tak et al. [10], utilizing spectro-temporal graph attention to capture time-frequency features. The final output is a two-dimensional logits score, which is optimized through weighted cross-entropy (WCE) loss.
3.2. WPT-SSL-AASIST
@@
(3) $L=\frac{1}{\sqrt{2}}[1,1]^{T},H=\frac{1}{\sqrt{2}}[1,-1]^{T}.$
@@
(4) $T_{LL},\left\{T_{LH},T_{HL},T_{HH}\right\}=\operatorname{DWT}(T).$
@@ 4.1. Dataset
Table 1. Statistics of all-type ADD benchmark in terms of training, development, and evaluation set.
@@
Speech-19LA. The 19LA dataset is one of the most popular datasets in the field of ADD. It contains a total of 12,456 real and 108,978 fake speeches generated using 11 Text-to-Speech (TTS) and 8 Voice Conversion (VC) spoofing algorithms (A01-A19). The real source domain of 19LA from VCTK [42]. The training and development sets consist of data from spoofing systems denoted by A01 to A06, while the evaluation set includes other generation methods denoted by A07 to A19. By testing on data where the generation method does not appear in the training dataset, we can evaluate the effectiveness of the CM.
@@
Sound-Codecfake-A3. Considering the principles of data selection and the total volume of the dataset, we chose the Codecfake A3 subset for sound. The real source domain is from the training subset of Audiocaps [43], and the fake sounds are generated using AudioGen based on the corresponding caption. Since some of the original Audiocaps audio links are no longer active, a small number of real sounds could not be downloaded. Ultimately, this condition includes 49,274 real sounds and 49,838 fake sounds. We randomly divided all the sound data into training, validation, and evaluation sets in a ratio of 7:1:2.
Singing voice-CtrSVDD. SVDD [4] is the first singing voice detection challenge. The real source domain of CtrSVDD includes multiple open-source Mandarin and Japanese singing datasets, and the synthesis methods encompass 14 different Singing Voice Synthesis (SVS) and Singing Voice Conversion (SVC) methods (A01-A14). The training protocol adheres to the principle that the generation methods used in the evaluation set are not visible in the training set, utilizing A01-A08 for training and A09-A14 for testing. We strictly follow the original training, validation, and evaluation split protocol of CtrSVDD.
@@
Music-FakeMusicCaps. FakeMusicCaps [9] is a deepfake music detection dataset. The real source domain of FakeMusicCaps is the MusicCaps [44] dataset, which consists of 5.5k 10-second music clips from AudioSet [45], each paired with an annotation by a professional musician. The fake-generated captions are identical to those in SunoCaps [46], and the synthesis methods include six approaches, labeled as TTM01-TTM05, along with one unknown method. We have restructured the dataset based on the fake methods. The real music is divided into a 7:1:2 ratio for training, validation, and testing. The training set for fake music includes TTM01-TTM03, the validation set includes TTM04, and the testing set includes TTM05 and the unknown method.
4.2. Baseline Countermeasure
@@
In our benchmark, we have five fundamental models, which we refer to as Spec-Resnet, AASIST, MERT-AASIST, WavLM-AASIST, and XLSR-AASIST based on their front-end and back-end concatenation.
Spec-Resnet is a classic audio classification method that combines Spectrogram with Resnet [47]. Although traditional features have previously underperformed compared to SSL features, further research is needed to study their generalizability across different types.
AASIST is a powerful deepfake CM that takes raw waveforms as input. It uses a sinc convolution layer [48] and six residual blocks to extract high-level representations, which are then fed into spectral and temporal graph attention for binary classification.
For SSL-AASIST, we employ three types of SSL features: MERT  [^3] https://huggingface.co/m-a-p/MERT-v1-330M, WavLM  [^4] https://huggingface.co/microsoft/wavlm-large, and XLSR  [^5] https://huggingface.co/facebook/wav2vec2-xls-r-300m. MERT is a self-supervised acoustic music understanding model that has achieved state-of-the-art (SOTA) performance in multiple music information retrieval (MIR) tasks. We believe it has the potential to distinguish deepfake music. When WavLM was proposed, it exhibited the SOTA performance among SSL models for all speech downstream tasks. Its performance in the ADD domain, including all-type ADD task, warrants further investigation. XLSR is the most representative of SSL for ADD tasks, and it has been proven to be the best SSL feature in the ADD domain across multiple SSL surveys [19, 49].
Regarding the training paradigms for SSL-AASIST, this paper explores four different approaches: FR-SSL-AASIST, FT-SSL-AASIST, PT-SSL-AASIST, and WPT-SSL-AASIST. Among these, FR and FT represent two distinct training paradigms: freezing and fine-tuning. In the FR method, all SSL parameters remain frozen, with only the final-layer SSL features fed into the AASIST back-end. FT method follows the same approach as FR but allows the SSL parameters to be updated through gradient backpropagation during training. The PT and WPT approach we have introduced in Section 3.
4.3. Training and Evaluation Protocol
@@
4.4. Implementation Details
For the pre-processing of the ADD baseline models, all audio samples were first down-sampled to 16,000 Hz and trimmed or padded to 64600 samples (same as the original AASIST and SSL-AASIST). For the Spec-Resnet, the spectrogram was computed with the number of FFT points set to 512, the hop length set to 160, and the window length set to 512. The back-end Resnet used Resnet18 followed by a fully connected layer to down-sample to 2 dimensions. For the training paradigm, FT-SSL-AASIST adopted the training parameters from Tak et al. [10], with an initial learning rate of $10^{-6}$ and a batch size of 14. For FR, PT, and WPT, an initial learning rate of $5^{-4}$ and a batch size of 32 were used. The dimensions of the 4-second audio processed by the SSL feature extractor are (201, 1024). For PT, the optimal number of prompt tokens was 10, while for WPT, the number of wavelet tokens was 4 and the number of regular prompt tokens was 6. In the single-type training experiments, the epoch was set to 50, and the learning rate was halved every 10 steps. In the co-training experiments, the epoch was set to 20, and the learning rate was halved every 4 steps.
5. Experiments
5.1. Investigation for Single-Type Training
Table 2. EER (%) results of the countermeasures (frozen SSL) trained on single-type ADD training set.
@@ 5.1. Investigation for Single-Type Training
In this section, we used a single-type training set to train deepfake CMs. First, we validated the performance of Spec-Resnet, AASIST, and FR-SSL-AASIST, as shown in Table 2. By analyzing the frozen SSL features, we can observe the inherent deepfake detection capabilities of SSL. Firstly, for the speech-trained CM, the best performance was achieved by XLSR-AASIST, which obtained the lowest in-domain EER of 1.28% and the lowest average EER of 32.58%. This result aligns with our expectations. Additionally, there is a significant drop in performance for cross-type ADD task. We observed that the singing test resulted in an EER of 29.72%, which is significantly better than the nearly 50% EER for sound and music. This denotes that there may be some shared characteristics in distinguishing between deepfake speech and deepfake singing. For the Sound task, AASIST achieved the lowest ID EER of 0.43%. However, its poor performance on speech and singing resulted in an average performance that was inferior to XLSR-AASIST. We believe that AASIST’s strong performance on sound is due to the fact that Codecfake-A3 uses only one deepfake method, leading to some degree of over-fitting. Nonetheless, its performance on music suggests that this over-fitting also benefits music detection. This indicates that there may be some commonality in distinguishing between sound and music, which are types of audio that usually do not contain human voices. For the singing task, XLSR-AASIST achieved both the best ID and best average performance. It obtained an ID EER of 9.45% and an average EER of 23.16%. Notably, XLSR-AASIST, which was trained solely on singing data, also demonstrated an EER of 12.89% on the speech task. This further indicates the commonality between speech and singing. For the music task, MERT-AASIST achieved the best ID and average performance with EER of 7.62% and 28.63%, respectively. This aligns with our expectations upon introducing MERT features in ADD task.
Table 3. EER (%) results of the countermeasures (finetuned SSL) trained on single-type ADD training set.
@@
Table 4. EER (%) comparison with different number of token.
@@
Table 5. EER (%) comparison with different paradigms.
@@
Figure 3. Different paradigms of PT-SSL-AASIST.
5.2. Prompt Tuning Countermeasures
@@
After determining the number of tokens, we investigated three paradigms for PT-SSL-AASIST, including Shallow-PT, After-PT, and Del-PT, as shown in Figure 3. Shallow-PT refers to inserting learnable prompts only in the first transformer encoder layer, which can demonstrate the importance of the deep paradigm where prompts are inserted in each layer. After-PT places the prompt position after the audio token, which might be effective due to the artifact information located in the silent region at the beginning of the audio [50]. Del-PT involves deleting the prompt token in the last layer, using only the audio tokens for classification, a method considered effective in some vision tasks [37]. Experimental results indicate that our proposed PT-SSL-AASIST paradigm is optimal, where prompts are inserted in each layer and the final layer combines the prompt and audio tokens for input into AASIST.
Table 6. EER (%) and training parameters comparison with different paradigms of speech-trained XLSR-AASIST.
@@
Figure 4. Convergence speed of different training paradigms.
WPT-SSL-AASIST. After deciding the PT architecture, we introduced WPT, applying DWT to the first four of the ten prompt tokens to better capture the frequency information of the audio. We compared the performance of FR, FT, PT, and WPT using the speech-trained XLSR-AASIST, as shown in Table 6. It can be observed that WPT achieved the best results compared to PT, obtaining a 0.15% EER on the ID speech evaluation set, with an average EER of 27.55%. Moreover, WPT does not increase the number of training parameters compared to PT, and compared to FT, the training parameters are reduced by 458 times. Overall, WPT outperformed FT, which in turn outperformed PT, and PT outperformed FR. We also recorded the training convergence speeds, as shown in Figure 4. It can be observed that FR, PT, and WPT converged significantly faster than FT, and both FT and PT showed less fluctuation compared to FR during convergence.
5.3. Co-trained Countermeasures
Table 7. EER (%) results for the countermeasures co-trained on the complete ADD training set.
@@ 5.3. Co-trained Countermeasures
Figure 5. T-SNE visualization for FT-XLSR-AASIST (left) and WPT-XLSR-AASIST (right). Different colors indicate features from different types: blue=speech, green=sound, orange=singing, purple=music. Different shapes represent different categories: cross=real, point=fake.
Figure 6. Attention map of the final transformer in co-trained XLSR-AASIST. Each column corresponds to the same deepfake audio sample. For both PT and WPT, we magnified the position of prompt tokens (1-10).
@@ 5.4. Interpretability
5.5. Conclusion
In this paper, we are dedicated to studying the all-type ADD task. We are the first to establish a comprehensive benchmark for evaluating the performance of current CMs on the all-type ADD task. Building on this foundation, we propose the PT-SSL training paradigm, which maintains performance while significantly reducing training parameters. Finally, to achieve all-type CM, we propose the WPT-SSL training paradigm, which leverages wavelet prompts to capture the type-invariant auditory deepfake information of SSL features. Our proposed co-trained WPT-XLSR-AASIST achieves an average EER of 3.58% across all-type ADD evaluation set.

### v2

@@
Abstract
The rapid advancement of audio generation technologies has escalated the risks of malicious deepfake audio across speech, sound, singing voice, and music, threatening multimedia security and trust. While existing countermeasures (CMs) perform well in single-type audio deepfake detection (ADD), their performance declines in cross-type scenarios. This paper is dedicated to studying the all-type ADD task. We are the first to comprehensively establish an all-type ADD benchmark to evaluate current CMs, incorporating cross-type deepfake detection across speech, sound, singing voice, and music. Then, we introduce the prompt tuning self-supervised learning (PT-SSL) training paradigm, which optimizes SSL front-end by learning specialized prompt tokens for ADD, requiring 458× fewer trainable parameters than fine-tuning (FT). Considering the auditory perception of different audio types, we propose the wavelet prompt tuning (WPT)-SSL method to capture type-invariant auditory deepfake information from the frequency domain without requiring additional training parameters, thereby enhancing performance over FT in the all-type ADD task. To achieve an universally CM, we utilize all types of deepfake audio for co-training. Experimental results demonstrate that WPT-XLSR-AASIST achieved the best performance, with an average EER of 3.58% across all evaluation sets.
Introduction
With the development of audio language model (ALM) technology, it has become increasingly easy to synthesize any type of audio, including deepfake speech, sound, singing voice, and music. These deepfake audios pose a threat to society in various fields such as media, entertainment, cybersecurity, and political communication. Fortunately, research on audio deepfake detection (ADD) has been increasing annually. Among these, the earliest studies focused on deepfake speech detection. Researchers have developed a series of deepfake countermeasures (CMs) aimed at effectively detecting deepfake speech, based on the ASVspoof challenges (Todisco et al. 2019; Liu et al. 2023; Wang et al. 2024). Currently, some ADD research has gone beyond speech, such as the detection of deepfake singing voices (Zhang et al. 2024; Xie et al. 2024a; Zang et al. 2024), sounds (Xie et al. 2024b, 2025), and music (Comanducci et al. 2024).
Figure 1: The challenge for current single-type trained CMs toward cross-type ADD task, highlighting the effectiveness of our proposed WPT-SSL CM.
@@ Introduction
For CMs, the most effective approach currently is to use pre-trained self-supervised learning (SSL) features along with a classification backbone. A representative CM in speech deepfake detection is XLSR-AASIST (Tak et al. 2022), which fine-tunes (FT) the wav2vec2-xls-r (XLSR) (Babu et al. 2022) model on speech deepfake detection dataset, achieving excellent intra-domain (ID) and out-of-domain (OOD) generalization performance. However, when dealing with the all-type ADD task, several challenges are encountered. Firstly, from the data perspective, it is uncertain whether a CM trained on single audio type can generalize to detect other types of deepfake audio. Although some studies have investigated cross-type detection for two types (Gohari et al. 2025; Xie et al. 2025), there has been no exploration of cross-type detection for all audio types. Secondly, there has been no investigation into whether a domain-invariant feature exists that can ensure the invariance of authenticity discrimination across different audio types. This requires a detailed investigation of various SSL features as well as handcrafted features. Lastly, concerning the algorithm, although fine-tuning can yield promising results, it is highly dependent on specific hyper-parameters and requires a significant amount of training parameters (Wang et al. 2025).
@@
To address the aforementioned challenges, in this paper, we aim to develop an all-type audio deepfake CM. We are the first to comprehensively establish an all-type ADD benchmark, which includes cross-type deepfake detection among speech, sound, singing voice, and music. For the feature of CMs, we investigate handcrafted features, raw waveforms, and various SSL-based features through both freezing and fine-tuning. For the back-end classifier, we use AASIST (Jung et al. 2022), the most popular model in the field of ADD, as the back-end, and combine it with SSL front-end to form SSL-AASIST.
@@
To efficiently optimize SSL front-end, inspired by Visual Prompt Tuning (VPT) (Jia et al. 2022), we proposed the Prompt Tuning (PT)-SSL training paradigm for ADD task. PT-SSL introduces learnable prompt tokens before the input of each transformer layer, while keeping the other parameters of the layers frozen, with the goal of learning specialized prompt tokens for the ADD task. Furthermore, considering the human perception of different audio types, the primary differences in perceiving audio types lie in their frequency domain distributions (Norman-Haignere et al. 2015; Kell et al. 2018; Munkong and Juang 2008). However, current SSL models like wav2vec2, which are primarily designed for speech recognition, focus on temporal and specific speech frequency information, lacking the ability to capture full-frequency information. To enhance frequency domain adaptability and enable SSL-based CM to quickly adapt to all types of deepfake audio, we propose wavelet prompt learning (WPT)-SSL method. WPT-SSL applies a discrete wavelet transform (DWT) to a portion of the prompt tokens, obtaining tokens for different frequency bands, thereby enhancing the full-frequency perception capability of SSL-based CM. Surprisingly, we discovered that WPT-SSL can learn a type-invariant deepfake detection prompt in a specific frequency band (HH) obtained through wavelet decomposition, thereby enabling all-type audio deepfake detection.
@@
Figure 2: Our proposed PT-SSL-AASIST (left) and WPT-SSL-AASIST (right). The differences between PT and WPT are illustrated below. WPT enhances the full-frequency perception of SSL-AASIST by applying DWT to part of the prompt tokens.
Prompt Tuning Countermeasure
@@
$Z_{i},E_{i}]={\color[rgb]{0,0,1}L_{i}}([{\color[rgb]{1,0,0}P_{i}},E_{i-1}]),\quad\text{for }i=2,3,...,l.$ (2)
Taking the most commonly used SSL feature in the ADD domain, XLSR-300m, as an example, after the final 24 layers, we obtain a matrix output $Z_{24},E_{24}]. II$ will serve as the input to AASIST. The back-end AASIST classifier fully follows the structure of SSL-AASIST by Tak et al. (Tak et al. 2022), utilizing spectro-temporal graph attention to capture time-frequency features. The final output is a two-dimensional logits score, which is optimized through weighted cross-entropy (WCE) loss.
WPT-SSL-AASIST
@@
$L=\frac{1}{\sqrt{2}}[1,1]^{T},H=\frac{1}{\sqrt{2}}[1,-1]^{T}.$ (3)
@@
$T_{LL},\left\{T_{LH},T_{HL},T_{HH}\right\}=\operatorname{DWT}(T).$ (4)
@@
Table 1: Statistics of all-type ADD benchmark in terms of training, development, and evaluation set.
Speech-19LA. A widely used benchmark containing 12,456 real and 108,978 fake samples from 11 TTS and 8 VC systems (A01–A19), with A01–A06 for training and A07–A19 for evaluation, ensuring no overlap in spoofing methods between sets.
@@
Sound-Codecfake-A3. We chose the Codecfake A3 subset for sound experiments. The real source domain is from the training subset of Audiocaps (Kim et al. 2019), and the fake sounds are generated using AudioGen based on the corresponding caption. This condition includes 49,274 real sounds and 49,838 fake sounds. We randomly divided all the sound data into training, validation, and evaluation sets in a ratio of 7:1:2.
Singing voice-CtrSVDD. SVDD (Zhang et al. 2024) is the first singing voice detection challenge. From the SVDD challenge, built on Mandarin and Japanese singing datasets with 14 SVS/SVC systems. A01–A08 are used for training, A09–A14 for testing, following the original protocol.
@@
Music-FakeMusicCaps. FakeMusicCaps (Comanducci et al. 2024) is a deepfake music detection dataset. The real source domain of FakeMusicCaps is the MusicCaps (Agostinelli et al. 2023) dataset, which consists of 5.5k 10-second music clips from AudioSet (Gemmeke et al. 2017), each paired with an annotation by a professional musician. Built from MusicCaps as real source and SunoCaps-style captions to synthesize fake music using six methods (TTM01–TTM05 + unknown). Real clips are split 7:1:2, while fake music uses TTM01–TTM03 for training, TTM04 for validation, and TTM05 + unknown for testing.
Baseline Countermeasure
@@
We establish five baseline models—Spec-Resnet, AASIST, and three SSL-enhanced variants: MERT-AASIST, WavLM-AASIST, and XLSR-AASIST—defined by their front-end and AASIST back-end combinations.
Spec-Resnet uses spectrograms with ResNet (He et al. 2016), representing traditional feature-based methods. Though often outperformed by SSL features, its cross-type generalization merits further study.
AASIST is a strong deepfake detection model operating directly on waveforms. It extracts high-level features via sinc convolutions (Ravanelli and Bengio 2018) and residual blocks, followed by spectral-temporal attention for binary classification.
SSL-AASIST variants incorporate:
  • MERT, an SSL model designed for music understanding with strong performance in MIR tasks;
  • WavLM, excelling in speech tasks and requiring further study in ADD scenarios;
  • XLSR, widely recognized as the most effective SSL feature for cross-lingual and cross-type ADD tasks (Phukan et al. 2024; Pascu et al. 2024).
We further explore four training paradigms for SSL-AASIST: FR (frozen) and FT (fine-tuned) differ by whether SSL parameters are updated during training; PT and WPT, our proposed methods, are described in Section Prompt Tuning Countermeasure.
Training and Evaluation Protocol
@@
Implementation Details
For the pre-processing of the ADD baseline models, all audio samples were first down-sampled to 16,000 Hz and trimmed or padded to 64600 samples (same as the original AASIST and SSL-AASIST). For the Spec-Resnet, the spectrogram was computed with the number of FFT points set to 512, the hop length set to 160, and the window length set to 512. The back-end Resnet used Resnet18 followed by a fully connected layer to down-sample to 2 dimensions. For the training paradigm, FT-SSL-AASIST adopted the training parameters from Tak et al. (Tak et al. 2022), with an initial learning rate of $10^{-6}$ and a batch size of 14. FR, PT, and WPT used a learning rate of $5\times 10^{-4}$ and batch size 32. SSL features had shape (201, 1024) for 4s audio. For single-type training, models trained for 50 epochs, halving the learning rate every 10 steps. Co-training ran for 20 epochs with LR halved every 4 steps.
Experiments
Investigation for Single-Type Training
@@
Table 2: EER (%) results of the countermeasures (frozen SSL) trained on single-type ADD training set.
In this section, we trained deepfake countermeasures (CMs) using single-type datasets and evaluated Spec-Resnet, AASIST, and FR-SSL-AASIST, as shown in Table 2. This setup allows us to assess the inherent detection capabilities of frozen SSL features. For speech, XLSR-AASIST performed best, achieving the lowest in-domain (ID) EER (1.28%) and average EER (32.58%). It also showed reasonable transferability to singing (EER: 29.72%)—much better than near-chance performance on sound and music—suggesting shared features between speech and singing. For sound, AASIST achieved the best in-domain EER (0.4%), likely due to overfitting on the single-method CodecFake-A3 data. Interestingly, this overfitting also benefitted music detection, hinting at similarities between non-speech audio types. For singing, XLSR-AASIST again led with an ID EER of 9.45% and average EER of 23.16%, and even generalized well to speech (EER: 12.89%), reinforcing the speech-singing similarity. For music, MERT-AASIST achieved the best performance (ID: 7.62%, Avg: 28.63%), consistent with MERT’s strength in music representation.
@@
Table 3: EER (%) results of the countermeasures (finetuned SSL) trained on single-type ADD training set.
@@
Table 4: EER (%) comparison with different number of token.
@@
Table 5: EER (%) comparison with different paradigms.
Figure 3: Different paradigms of PT-SSL-AASIST.
Prompt Tuning Countermeasures
@@
After determining the number of tokens, we investigated three paradigms for PT-SSL-AASIST, including Shallow-PT, After-PT, and Del-PT, as shown in Figure 3. Shallow-PT refers to inserting learnable prompts only in the first transformer encoder layer, which can demonstrate the importance of the deep paradigm where prompts are inserted in each layer. After-PT places the prompt position after the audio token, which might be effective due to the artifact information located in the silent region at the beginning of the audio (Zhang et al. 2023). Del-PT involves deleting the prompt token in the last layer, using only the audio tokens for classification, a method considered effective in some vision tasks (Darcet et al. 2024). Experimental results indicate that our proposed PT-SSL-AASIST paradigm is optimal, where prompts are inserted in each layer and the final layer combines the prompt and audio tokens for input into AASIST.
WPTs PTs Speech Sound Singing Music AVG
0 10 0.22 47.26 33.84 41.85 30.79
2 8 0.16 49.18 38.22 34.84 30.60
4 6 0.15 45.36 33.32 28.61 26.86
6 4 0.16 47.86 36.21 31.52 28.94
8 2 0.18 47.40 40.68 32.25 30.13
10 0 0.11 49.40 36.97 43.68 32.54
Table 6: EER (%) comparison with different number of wavelet prompt tokens (WPTs) and prompt tokens (PTs).
@@
Table 7: EER (%) and training parameters comparison with different paradigms of speech-trained XLSR-AASIST.
Figure 4: Convergence speed of different paradigms.
@@
Table 8: EER (%) results for the countermeasures co-trained on the complete ADD training set.
WPT-SSL-AASIST. After deciding the PT architecture, we introduced WPT, applying DWT to a part of the ten prompt tokens to better capture the frequency information of the audio. We first investigate the optimal number of tokens by comparing different settings of wavelet prompt tokens (WPTs) and standard prompt tokens (PTs). As shown in Table 6, the best overall performance is achieved when WPTs = 4, yielding an average EER of 26.86%. Interestingly, this configuration also naturally aligns each of the four frequency bands (LL, LH, HL, HH) with a single token. Although the configuration with WPTs = 10 achieves the lowest EER of 0.11% on the speech type, its performance degrades significantly on other types, making it less favorable overall. Based on the optimal performance observed, we adopt 4 WPTs for the remainder of our experiments.
Figure 5: T-SNE visualization for FT-XLSR-AASIST (left) and WPT-XLSR-AASIST (right). Different colors indicate features from different types: blue=speech, green=sound, orange=singing, purple=music. Different shapes represent different categories: cross=real, point=fake.
Figure 6: Attention map of the final transformer in co-trained XLSR-AASIST. Each column corresponds to the same deepfake audio sample. For both PT and WPT, we magnified the position of prompt tokens (1-10).
Then, we compared the performance of FR, FT, PT, and WPT using the speech-trained XLSR-AASIST, as shown in Table 7. It can be observed that WPT achieved the best results compared to PT, obtaining a 0.15% EER on the ID speech evaluation set, with an average EER of 27.55%. Moreover, WPT does not increase the number of training parameters compared to PT, and compared to FT, the training parameters are reduced by 458 times. Overall, WPT outperformed FT, which in turn outperformed PT, and PT outperformed FR. We also recorded the training convergence speeds, as shown in Figure 4. It can be observed that FR, PT, and WPT converged significantly faster than FT, and both FT and PT showed less fluctuation compared to FR during convergence.
Co-trained Countermeasures
@@
Conclusion
In this paper, we are dedicated to studying the all-type ADD task. We are the first to establish a comprehensive benchmark for evaluating the performance of current CMs on the all-type ADD task. To achieve all-type CM, we propose the WPT-SSL training paradigm, which leverages wavelet prompts to capture the type-invariant auditory deepfake information of SSL features. Our proposed co-trained WPT-XLSR-AASIST achieves an average EER of 3.58% across all-type ADD evaluation set.
Acknowledgements
This work is supported by the Beijing Natural Science Foundation (No. L252143), and the National Natural Science Foundation of China (NSFC) (No.62201571, No.62101553, No.U21B20210).